City Index City Index City A 106 City F 144 City B 436 City G 190 City C 93 City H 100 City D 98 caty J 301 City E 74 City K99 The following table shows a housing index that can house valued at $210,000 in City H. Find the price of a comparable house in City A City E, and City B be used to compare housing prices in different cities. Suppose you seea If you know the price of a particular home in your town, you can use the index to find the price of a comparable house in another town with the formula below price price index (other town) other town) (your town) Index (your town) A comparable house in City A would cost about $ (Round to the nearest dollar) A comparable house in City E would cost about s Round to the nearest dollar) A comparable house in City B would cost about s (Round to the nearest dollar)
Answer #1

Solution:

Comparable house in City A = House price City H * Index (City A) / Index (City H) = 210,000 * 106/100 = $222,600

Comparable house in City E = House price City H * Index (City E) / Index (City H) = 210,000 * 74/100 = $155,400

Comparable house in City B = House price City H * Index (City B) / Index (City H) = 210,000 * 436/100 = $915,600